40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Jul 23 15:00:28 2017, Jamie, Omnistructure, VAC, strange sound around X arm vacuum pumps 
    Reply  Mon Jul 24 09:55:29 2017, Steve, Omnistructure, VAC, controller failer TP2controllerFails.png
       Reply  Tue Jul 25 08:48:57 2017, Steve, Update, VAC, RGA scan at d278 RGA278d.png
          Reply  Mon Aug 28 10:28:07 2017, Steve, Update, VAC, RGA scan 20 day after 17 Torr 20d_after_17_torr.png
             Reply  Tue Jan 30 10:22:20 2018, Steve, Update, VAC, RGA scan at day 175 RGApd80-b-468-175d.pngRgaOn20hrs.png
                Reply  Fri Mar 16 09:36:56 2018, Steve, Update, VAC, RGA scan at day 511,218d RgaScan511d.png08.png
                   Reply  Mon Mar 26 09:25:18 2018, Steve, Update, VAC, VM1 opened CC4VM1.png
                      Reply  Tue Mar 27 08:58:21 2018, Steve, Update, VAC, VM1 opened.......scan fine rga2018march27.png
          Reply  Tue Sep 19 11:05:26 2017, Steve, Update, VAC, RGA scan at day 334  RGAscan334d.png
             Reply  Fri Oct 27 10:14:11 2017, Steve, Update, VAC, RGA scan at day 372 RGA_scan_d372.png
                Reply  Mon Oct 30 10:09:41 2017, Steve, Update, VAC, TP2 failed tp2failed.pngifo_1.0E-5_Torrit.pngtp2failed2dago.png4days.png
                   Reply  Mon Nov 6 09:09:43 2017, Steve, Update, VAC, setting up new TP2 turbo 
       Reply  Tue Jul 25 14:04:06 2017, Steve, Update, VAC, turbo controller installed and we are running at vac normal TP2@75Krpm.png
Message ID: 13142     Entry time: Tue Jul 25 08:48:57 2017     In reply to: 13134     Reply to this: 13260   13321
Author: Steve 
Type: Update 
Category: VAC 
Subject: RGA scan at d278 

The RGA did not shut down at the turbo pump controller failing.

Quote:

Ifo pressure was 5.5 mTorr this morning. The PSL shutter was still open. TP2 controller failed. Interlock closed V1, V4 and VM1

Turbo pump 2 is the fore pump of the Maglev. The pressure here was 3.9 Torr so The Magelv got warm ~38C but it was still rotating at 560 Hz normal with closed V1

What I did:

Looked at pressures of Hornet and Super Bee  Instru Tech. Inc

Closed all annuloses and VA6,  disconnected V4 and VA6 and turned on external fan to cool Maglev

Opened V7 to pump the Maglev fore line with TP3

V1 opened manually when foreline pressure dropped to <2mTorr at P2 and the body temp of the Maglev cooled down to  25-27 C

VM1 opened at 1e-5 Torr

Valve configuration: vacuum normal with annuloses not pumped

Ifo pressure 8.5e-6 Torr -IT at 10am,  P2 foreline pressure 64 mTorr, TP3 controller 0.17A   22C  50Krpm

note: all valves open manually, interlock can only close them

 

Quote:

While walking down to the X end to reset c1iscex I heard what I would call a "rythmic squnching" sound coming from under the turbo pump.  I would have said the sound was coming from a roughing pump, but none of them are on (as far as I can tell).

Steve maybe look into this??

PS: please call me next time you see the vacuum is not Vacuum Normal

 

Attachment 1: RGA278d.png  40 kB  Uploaded Tue Jul 25 09:49:20 2017  | Hide | Hide all
RGA278d.png
ELOG V3.1.3-